# Model-Agnostic Bounds for Augmented Inverse Probability Weighted Estimators’ Wald-Confidence Interval Coverage in Randomized Controlled Trials

### v1

@@ 6 Discussion
$\quad+{\mathcal{C}}\frac{\varrho_{\#,a}}{\varsigma_{\#,a}^{3}\sqrt{n}}+2\exp\left\{-\frac{\eta_{7}^{2}n}{2\varsigma_{\#,a}^{2}}\right\}+t_{2}^{-q_{2}/2}+{\mathbbm{1}}(\psi_{*,a}\neq 0)t_{4}^{-q_{4}/2}+t_{6}^{-q_{6}/2}+\frac{16M^{2}(1-\tau_{\pi})^{2}{\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}}{\tau_{\pi}^{2}\eta_{8}^{2}}.$
Additionally under Condition S2, the last term on the right-hand side can be replaced by
$2\exp\left\{-\frac{\tau_{\pi}^{2}\eta_{8}^{2}}{2\tilde{s}_{3}^{2}M^{2}(1-\tau_{\pi})^{2}{\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}}\right\}.$
@@
${\mathbb{E}}[\tilde{\sigma}_{a}^{2}]$ $=n^{-1/2}{\mathbb{E}}[{\mathbb{G}}_{n}\{{\mathcal{T}}_{a}(\tilde{Q}_{a})^{2}-{\mathcal{T}}_{a}(Q_{\#,a})^{2}\}]+{\mathbb{E}}[P_{*}{\mathcal{T}}_{a}(\tilde{Q}_{a})^{2}]-{\mathbb{E}}[\tilde{\psi}_{a}^{2}].$ (S8)
@@
$\Pr\left(|\tilde{\sigma}_{a}^{2}-{\mathbb{E}}[\tilde{\sigma}_{a}^{2}]|>\sum_{j=1}^{8}\eta_{j},E\right)$
$\leq\Pr\left(n^{-1/2}\left|{\mathbb{G}}\{{\mathcal{T}}_{a}(\tilde{Q}_{a})^{2}-{\mathcal{T}}_{a}(Q_{\#,a})^{2}\}-{\mathbb{E}}[{\mathbb{G}}\{{\mathcal{T}}_{a}(\tilde{Q}_{a})^{2}-{\mathcal{T}}_{a}(Q_{\#,a})^{2}\}]\right|>\eta_{6},E\right)$
$\quad+\Pr(|(P_{n}-P_{*}){\mathcal{T}}_{a}(Q_{\#,a})^{2}|>\eta_{7})+\Pr(|P_{*}{\mathcal{T}}_{a}(\tilde{Q}_{a})^{2}-{\mathbb{E}}[P_{*}{\mathcal{T}}_{a}(\tilde{Q}_{a})^{2}]|>\eta_{8})$
$\quad+\Pr\left(\left|\tilde{\psi}_{a}^{2}-{\mathbb{E}}[\tilde{\psi}_{a}^{2}]\right|>\sum_{j=1}^{5}\eta_{j}\right).$
@@
$\sqrt{n}\eta_{6}$ $\geq|{\mathbb{E}}[{\mathbb{G}}\{{\mathcal{T}}_{a}(\tilde{Q}_{a})^{2}-{\mathcal{T}}_{a}(Q_{\#,a})^{2}\}]|+{\mathcal{C}}(1+\alpha){\mathbb{E}}\|{\mathbb{G}}_{n}\|_{{\mathcal{F}}^{{\mathcal{T}}_{a}^{2}}_{\delta F^{{\mathcal{T}}_{a}^{2}}}}+{\mathcal{C}}q\left\{\left(\delta F^{{\mathcal{T}}_{a}^{2}}+\frac{F^{{\mathcal{T}}_{a}^{2}}}{\sqrt{n}}\right)\sqrt{t}+\frac{F^{{\mathcal{T}}_{a}^{2}}t}{\alpha\sqrt{n}}\right\}$
$>|{\mathbb{E}}[{\mathbb{G}}\{{\mathcal{T}}_{a}(\tilde{Q}_{a})^{2}-{\mathcal{T}}_{a}(Q_{\#,a})^{2}\}]|,$
@@
$\Pr\left(n^{-1/2}\left|{\mathbb{G}}\{{\mathcal{T}}_{a}(\tilde{Q}_{a})^{2}-{\mathcal{T}}_{a}(Q_{\#,a})^{2}\}-{\mathbb{E}}[{\mathbb{G}}\{{\mathcal{T}}_{a}(\tilde{Q}_{a})^{2}-{\mathcal{T}}_{a}(Q_{\#,a})^{2}\}]\right|>\eta_{6},E\right)\leq t_{6}^{-q_{6}/2}.$
@@
$\Pr(|(P_{n}-P_{*}){\mathcal{T}}_{a}(Q_{\#,a})^{2}|>\eta_{7})$ $=\Pr\left(\sqrt{n}|(P_{n}-P_{*}){\mathcal{T}}_{a}(Q_{\#,a})^{2}|/\varsigma_{\#,a}>\eta_{7}\sqrt{n}/\varsigma_{\#,a}\right)$
@@
$\Pr(|P_{*}{\mathcal{T}}_{a}(\tilde{Q}_{a})^{2}-{\mathbb{E}}[P_{*}{\mathcal{T}}_{a}(\tilde{Q}_{a})^{2}]|>\eta_{8})\leq\frac{16M^{2}(1-\tau_{\pi})^{2}{\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}}{\tau_{\pi}^{2}\eta_{8}^{2}}.$
Additionally under Condition S2,
$\Pr(|P_{*}{\mathcal{T}}_{a}(\tilde{Q}_{a})^{2}-{\mathbb{E}}[P_{*}{\mathcal{T}}_{a}(\tilde{Q}_{a})^{2}]|>\eta_{8})\leq 2\exp\left\{-\frac{\tau_{\pi}^{2}\eta_{8}^{2}}{2\tilde{s}_{3}^{2}M^{2}(1-\tau_{\pi})^{2}{\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}}\right\}.$
Finally, $\Pr\left(\left|\tilde{\psi}_{a}^{2}-{\mathbb{E}}[\tilde{\psi}_{a}^{2}]\right|>\sum_{j=1}^{5}\eta_{j}\right)$ can be bounded as in Lemma S13. ∎
@@
$\sigma_{\dagger,a}^{2}-\sigma_{\#,a}^{2}=n^{-1/2}{\mathbb{E}}[{\mathbb{G}}_{n}\{{\mathcal{T}}_{a}(\tilde{Q}_{a})^{2}-{\mathcal{T}}_{a}(Q_{\#,a})^{2}\}]+P_{*}\left\{{\mathbb{E}}_{\tilde{Q}_{a}}[{\mathcal{T}}_{a}(\tilde{Q}_{a})^{2}]-{\mathcal{T}}_{a}(Q_{\#,a})^{2}\right\}-{\mathrm{Var}}(\tilde{\psi}_{a}).$
@@
$P_{*}\left\{{\mathbb{E}}_{\tilde{Q}_{a}}[{\mathcal{T}}_{a}(\tilde{Q}_{a})^{2}]-{\mathcal{T}}_{a}(Q_{\#,a})^{2}\right\}$
$=P_{*}\left\{\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}{\mathbb{E}}_{\tilde{Q}_{a}}[(\tilde{Q}_{a}-Q_{\#,a})^{2}]\right\}+2P_{*}\left\{\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}(Q_{\#,a}-Q_{*,a}){\mathbb{E}}_{\tilde{Q}_{a}}[\tilde{Q}_{a}-Q_{\#,a}]\right\}.$
@@
$|{\mathbb{E}}[{\mathbb{G}}_{n}\{{\mathcal{T}}_{a}(\tilde{Q}_{a})^{2}-{\mathcal{T}}_{a}(Q_{\#,a})^{2}\}]|$
$={\mathbb{E}}[{\mathbb{G}}_{n}\{{\mathcal{T}}_{a}(\tilde{Q}_{a})^{2}-{\mathcal{T}}_{a}(Q_{\#,a})^{2}\}{\mathbbm{1}}(\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}\leq 2\delta M)]$
$\qquad+{\mathbb{E}}[{\mathbb{G}}_{n}\{{\mathcal{T}}_{a}(\tilde{Q}_{a})^{2}-{\mathcal{T}}_{a}(Q_{\#,a})^{2}\}{\mathbbm{1}}(\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}>2\delta M)]$
$\leq{\mathbb{E}}\|{\mathbb{G}}_{n}\|_{{\mathcal{F}}^{{\mathcal{T}}_{a}^{2}}_{\delta F^{{\mathcal{T}}_{a}^{2}}}}+{\mathbb{E}}\|{\mathbb{G}}_{n}\|_{{\mathcal{F}}^{{\mathcal{T}}_{a}^{2}}}\Pr(\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}>2\delta M)$
$\leq{\mathcal{C}}F^{{\mathcal{T}}_{a}^{2}}J(2\delta,{\mathcal{F}},M)+{\mathcal{C}}\frac{F^{{\mathcal{T}}_{a}^{2}}J^{2}(2\delta,{\mathcal{F}},M)}{\delta^{2}\sqrt{n}}+{\mathcal{C}}F^{{\mathcal{T}}_{a}^{2}}J(2,{\mathcal{F}},M)\Pr(\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}>2\delta M),$ (S9)
where $\Pr(\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}>2\delta M)$ can be bounded by Markov’s inequality, or with subgaussian tail bound under Condition 10. Thus, (S4) holds. By the triangle inequality and Lemma S13,
@@
$\leq\frac{{\mathcal{C}}}{\sqrt{n}}F^{{\mathcal{T}}_{a}^{2}}J(2\delta,{\mathcal{F}},M)+{\mathcal{C}}\frac{F^{{\mathcal{T}}_{a}^{2}}J^{2}(2\delta,{\mathcal{F}},M)}{\delta^{2}n}+\frac{{\mathcal{C}}}{\sqrt{n}}F^{{\mathcal{T}}_{a}^{2}}J(2,{\mathcal{F}},M)\Pr(\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}>2\delta M)$
$\quad+P_{*}\left\{\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}{\mathbb{E}}_{\tilde{Q}_{a}}[(\tilde{Q}_{a}-Q_{\#,a})^{2}]\right\}+2\left|P_{*}\left\{\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}(Q_{\#,a}-Q_{*,a}){\mathbb{E}}_{\tilde{Q}_{a}}[\tilde{Q}_{a}-Q_{\#,a}]\right\}\right|$
@@
$\left|\Pr(\sqrt{n}(\tilde{\psi}_{a}-\psi_{*,a})\leq z_{\alpha}\tilde{\sigma}_{a})-(1-\alpha)-\phi(z_{\alpha})z_{\alpha}\frac{\sigma_{\dagger,a}-\sigma_{\#,a}}{\sigma_{\#,a}}\right|$
@@
$\quad+\Pr(|\sqrt{n}\{\tilde{\psi}_{a}-\psi_{*,a}-P_{n}D_{a}(Q_{\#,a},\psi_{*,a})\}|>\tilde{\eta},\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}\leq 2\delta M)$
$\quad+\Pr\left(|z_{\alpha}(\tilde{\sigma}_{a}-\sigma_{\dagger,a})|>\sum_{j=1}^{8}\tilde{\eta}_{j},\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}\leq 2\delta M\right)+\Pr(\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}>2\delta M).$
@@
$\Pr(|\sqrt{n}\{\tilde{\psi}_{a}-\psi_{*,a}-P_{n}D_{a}(Q_{\#,a},\psi_{*,a})\}|>\tilde{\eta},\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}\leq 2\delta M)$
$\leq\Pr(|{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})|>\tilde{\eta},\|H_{a}(\tilde{Q}_{a})\|_{2}\leq\delta F^{H_{a}})\leq\tilde{t}^{-\tilde{q}/2}.$
Next bound $\Pr\left(|z_{\alpha}(\tilde{\sigma}_{a}-\sigma_{\dagger,a})|>\sum_{j=1}^{8}\tilde{\eta}_{j},\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}\leq 2\delta M\right)$. This term is zero if $z_{\alpha}=0$ and $\sum_{j=1}^{8}\tilde{\eta}_{j}>0$. Otherwise, since $\tilde{\sigma}_{a}\geq 0$ and $\sigma_{\dagger,a}>0$, it holds that
$\Pr\left(|z_{\alpha}(\tilde{\sigma}_{a}-\sigma_{\dagger,a})|>\sum_{j=1}^{8}\tilde{\eta}_{j},\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}\leq 2\delta M\right)$
$\leq\Pr\left(|\tilde{\sigma}_{a}^{2}-\sigma_{\dagger,a}^{2}|>\sigma_{\dagger,a}\sum_{j=1}^{8}\tilde{\eta}_{j}/|z_{\alpha}|,\|{\mathcal{T}}_{a}(\tilde{Q}_{a})^{2}-{\mathcal{T}}_{a}(Q_{\#,a})^{2}\|_{2}\leq\delta F^{{\mathcal{T}}_{a}^{2}}\right).$
@@
$\Pr(\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}>2\delta M)\leq\frac{{\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}}{4\delta^{2}M^{2}}.$ (S10)
Additionally under Condition 10, if $2\delta M>{\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}$, then
$\Pr(\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}>2\delta M)\leq\exp\left\{-\frac{(2\delta M-{\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2})^{2}}{2\tilde{s}_{1}^{2}{\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}}\right\}.$ (S11)
@@
$\left|\Pr(\sqrt{n}(\tilde{\psi}_{a}-\psi_{*,a})\leq z_{\alpha}\tilde{\sigma}_{a})-(1-\alpha)-\phi(z_{\alpha})z_{\alpha}\frac{\sigma_{\dagger,a}-\sigma_{\#,a}}{\sigma_{\#,a}}\right|$
@@
$\quad+{\mathcal{C}}\frac{\varrho_{\#,a}}{\varsigma_{\#,a}^{3}\sqrt{n}}+2\exp\left\{-\frac{\eta_{7}^{2}n}{2\varsigma_{\#,a}^{2}}\right\}+t_{2}^{-q_{2}/2}+{\mathbbm{1}}(\psi_{*,a}\neq 0)t_{4}^{-q_{4}/2}+t_{6}^{-q_{6}/2}+\frac{16M^{2}(1-\tau_{\pi})^{2}{\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}}{\tau_{\pi}^{2}\eta_{8}^{2}}\Bigg]$
$\quad+\frac{{\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}}{4\delta^{2}M^{2}},$ (S12)
@@
$\frac{16M^{2}(1-\tau_{\pi})^{2}{\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}}{\tau_{\pi}^{2}\eta_{8}^{2}}\text{ can be replaced by }2\exp\left\{-\frac{\tau_{\pi}^{2}\eta_{8}^{2}}{2\tilde{s}_{3}^{2}M^{2}(1-\tau_{\pi})^{2}{\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}}\right\}$
under Condition S2, and
$\frac{{\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}}{4\delta^{2}M^{2}}\text{ can be replaced by }\exp\left\{-\frac{(2\delta M-{\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2})^{2}}{2\tilde{s}_{1}^{2}{\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}}\right\}$
if $2\delta M>{\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}$ under Condition 10.
@@
$J(2\delta,{\mathcal{F}},M)$ $\leq\int_{0}^{2\delta}\sqrt{C_{1}(\nu)+2(\nu-1)\log(1/\epsilon)}{\mathrm{d}}\epsilon$

### v2

@@ 6 Discussion
$\quad+{\mathcal{C}}\frac{\varrho_{\#,a}}{\varsigma_{\#,a}^{3}\sqrt{n}}+2\exp\left\{-\frac{\eta_{7}^{2}n}{2\varsigma_{\#,a}^{2}}\right\}+t_{2}^{-q_{2}/2}+{\mathbbm{1}}(\psi_{*,a}\neq 0)t_{4}^{-q_{4}/2}+t_{6}^{-q_{6}/2}+\frac{16M^{2}(1-\tau_{\pi})^{2}{E}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}}{\tau_{\pi}^{2}\eta_{8}^{2}}.$
Additionally under Condition 12, the last term on the right-hand side can be replaced by
$\tilde{a}_{2}n^{\tilde{d}_{2}}\exp\left\{-\tilde{c}_{2}\left(\frac{\eta_{8}}{\tilde{r}_{2}(n)}\right)^{\tilde{q}_{2}}\right\}+\tilde{b}_{2}n^{-1/2}\log n.$
@@
${E}[\tilde{\sigma}_{a}^{2}]$ $=n^{-1/2}{E}[{\mathbb{G}}_{n}\{{\mathcal{T}}_{a}(\tilde{Q}_{a})^{2}-{\mathcal{T}}_{a}(Q_{\#,a})^{2}\}]+{E}[P_{*}{\mathcal{T}}_{a}(\tilde{Q}_{a})^{2}]-{E}[\tilde{\psi}_{a}^{2}].$ (S8)
@@
${\mathrm{pr}}\left(|\tilde{\sigma}_{a}^{2}-{E}[\tilde{\sigma}_{a}^{2}]|>\sum_{j=1}^{8}\eta_{j},E\right)$
$\leq{\mathrm{pr}}\left(n^{-1/2}\left|{\mathbb{G}}\{{\mathcal{T}}_{a}(\tilde{Q}_{a})^{2}-{\mathcal{T}}_{a}(Q_{\#,a})^{2}\}-{E}[{\mathbb{G}}\{{\mathcal{T}}_{a}(\tilde{Q}_{a})^{2}-{\mathcal{T}}_{a}(Q_{\#,a})^{2}\}]\right|>\eta_{6},E\right)$
$\quad+{\mathrm{pr}}(|(P_{n}-P_{*}){\mathcal{T}}_{a}(Q_{\#,a})^{2}|>\eta_{7})+{\mathrm{pr}}(|P_{*}{\mathcal{T}}_{a}(\tilde{Q}_{a})^{2}-{E}[P_{*}{\mathcal{T}}_{a}(\tilde{Q}_{a})^{2}]|>\eta_{8})$
$\quad+{\mathrm{pr}}\left(\left|\tilde{\psi}_{a}^{2}-{E}[\tilde{\psi}_{a}^{2}]\right|>\sum_{j=1}^{5}\eta_{j}\right).$
@@
$\sqrt{n}\eta_{6}$ $\geq|{E}[{\mathbb{G}}\{{\mathcal{T}}_{a}(\tilde{Q}_{a})^{2}-{\mathcal{T}}_{a}(Q_{\#,a})^{2}\}]|+{\mathcal{C}}(1+\alpha){E}\|{\mathbb{G}}_{n}\|_{{\mathcal{F}}^{{\mathcal{T}}_{a}^{2}}_{\delta F^{{\mathcal{T}}_{a}^{2}}}}+{\mathcal{C}}q\left\{\left(\delta F^{{\mathcal{T}}_{a}^{2}}+\frac{F^{{\mathcal{T}}_{a}^{2}}}{\sqrt{n}}\right)\sqrt{t}+\frac{F^{{\mathcal{T}}_{a}^{2}}t}{\alpha\sqrt{n}}\right\}$
$>|{E}[{\mathbb{G}}\{{\mathcal{T}}_{a}(\tilde{Q}_{a})^{2}-{\mathcal{T}}_{a}(Q_{\#,a})^{2}\}]|,$
@@
${\mathrm{pr}}\left(n^{-1/2}\left|{\mathbb{G}}\{{\mathcal{T}}_{a}(\tilde{Q}_{a})^{2}-{\mathcal{T}}_{a}(Q_{\#,a})^{2}\}-{E}[{\mathbb{G}}\{{\mathcal{T}}_{a}(\tilde{Q}_{a})^{2}-{\mathcal{T}}_{a}(Q_{\#,a})^{2}\}]\right|>\eta_{6},E\right)\leq t_{6}^{-q_{6}/2}.$
@@
${\mathrm{pr}}(|(P_{n}-P_{*}){\mathcal{T}}_{a}(Q_{\#,a})^{2}|>\eta_{7})$ $={\mathrm{pr}}\left(\sqrt{n}|(P_{n}-P_{*}){\mathcal{T}}_{a}(Q_{\#,a})^{2}|/\varsigma_{\#,a}>\eta_{7}\sqrt{n}/\varsigma_{\#,a}\right)$
@@
${\mathrm{pr}}(|P_{*}{\mathcal{T}}_{a}(\tilde{Q}_{a})^{2}-{E}[P_{*}{\mathcal{T}}_{a}(\tilde{Q}_{a})^{2}]|>\eta_{8})\leq\frac{16M^{2}(1-\tau_{\pi})^{2}{E}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}}{\tau_{\pi}^{2}\eta_{8}^{2}}.$
Additionally under Condition 12,
${\mathrm{pr}}(|P_{*}{\mathcal{T}}_{a}(\tilde{Q}_{a})^{2}-{E}[P_{*}{\mathcal{T}}_{a}(\tilde{Q}_{a})^{2}]|>\eta_{8})\leq\tilde{a}_{2}n^{\tilde{d}_{2}}\exp\left\{-\tilde{c}_{2}\left(\frac{\eta_{8}}{\tilde{r}_{2}(n)}\right)^{\tilde{q}_{2}}\right\}+\tilde{b}_{2}n^{-1/2}\log n.$
Finally, ${\mathrm{pr}}\left(\left|\tilde{\psi}_{a}^{2}-{E}[\tilde{\psi}_{a}^{2}]\right|>\sum_{j=1}^{5}\eta_{j}\right)$ can be bounded as in Lemma S13. ∎
@@
$\sigma_{\dagger,a}^{2}-\sigma_{\#,a}^{2}=n^{-1/2}{E}[{\mathbb{G}}_{n}\{{\mathcal{T}}_{a}(\tilde{Q}_{a})^{2}-{\mathcal{T}}_{a}(Q_{\#,a})^{2}\}]+P_{*}\left\{{E}_{\tilde{Q}_{a}}[{\mathcal{T}}_{a}(\tilde{Q}_{a})^{2}]-{\mathcal{T}}_{a}(Q_{\#,a})^{2}\right\}-{\mathrm{var}}(\tilde{\psi}_{a}).$
@@
$P_{*}\left\{{E}_{\tilde{Q}_{a}}[{\mathcal{T}}_{a}(\tilde{Q}_{a})^{2}]-{\mathcal{T}}_{a}(Q_{\#,a})^{2}\right\}$
$=P_{*}\left\{\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}{E}_{\tilde{Q}_{a}}[(\tilde{Q}_{a}-Q_{\#,a})^{2}]\right\}+2P_{*}\left\{\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}(Q_{\#,a}-Q_{*,a}){E}_{\tilde{Q}_{a}}[\tilde{Q}_{a}-Q_{\#,a}]\right\}.$
@@
$|{E}[{\mathbb{G}}_{n}\{{\mathcal{T}}_{a}(\tilde{Q}_{a})^{2}-{\mathcal{T}}_{a}(Q_{\#,a})^{2}\}]|$
$={E}[{\mathbb{G}}_{n}\{{\mathcal{T}}_{a}(\tilde{Q}_{a})^{2}-{\mathcal{T}}_{a}(Q_{\#,a})^{2}\}{\mathbbm{1}}(\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}\leq 2\delta M)]$
$\qquad+{E}[{\mathbb{G}}_{n}\{{\mathcal{T}}_{a}(\tilde{Q}_{a})^{2}-{\mathcal{T}}_{a}(Q_{\#,a})^{2}\}{\mathbbm{1}}(\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}>2\delta M)]$
$\leq{E}\|{\mathbb{G}}_{n}\|_{{\mathcal{F}}^{{\mathcal{T}}_{a}^{2}}_{\delta F^{{\mathcal{T}}_{a}^{2}}}}+{E}\|{\mathbb{G}}_{n}\|_{{\mathcal{F}}^{{\mathcal{T}}_{a}^{2}}}{\mathrm{pr}}(\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}>2\delta M)$
$\leq{\mathcal{C}}F^{{\mathcal{T}}_{a}^{2}}J(2\delta,{\mathcal{F}},M)+{\mathcal{C}}\frac{F^{{\mathcal{T}}_{a}^{2}}J^{2}(2\delta,{\mathcal{F}},M)}{\delta^{2}\sqrt{n}}+{\mathcal{C}}F^{{\mathcal{T}}_{a}^{2}}J(2,{\mathcal{F}},M){\mathrm{pr}}(\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}>2\delta M),$ (S9)
where ${\mathrm{pr}}(\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}>2\delta M)$ can be bounded by Markov’s inequality, or with a sub-Weibull tail bound under Condition 11. Thus, (S4) holds. By the triangle inequality and Lemma S13,
@@
$\leq\frac{{\mathcal{C}}}{\sqrt{n}}F^{{\mathcal{T}}_{a}^{2}}J(2\delta,{\mathcal{F}},M)+{\mathcal{C}}\frac{F^{{\mathcal{T}}_{a}^{2}}J^{2}(2\delta,{\mathcal{F}},M)}{\delta^{2}n}+\frac{{\mathcal{C}}}{\sqrt{n}}F^{{\mathcal{T}}_{a}^{2}}J(2,{\mathcal{F}},M){\mathrm{pr}}(\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}>2\delta M)$
$\quad+P_{*}\left\{\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}{E}_{\tilde{Q}_{a}}[(\tilde{Q}_{a}-Q_{\#,a})^{2}]\right\}+2\left|P_{*}\left\{\frac{1-\pi_{*}(a\mid\cdot)}{\pi_{*}(a\mid\cdot)}(Q_{\#,a}-Q_{*,a}){E}_{\tilde{Q}_{a}}[\tilde{Q}_{a}-Q_{\#,a}]\right\}\right|$
@@
$\left|{\mathrm{pr}}(\sqrt{n}(\tilde{\psi}_{a}-\psi_{*,a})\leq z_{\alpha}\tilde{\sigma}_{a})-(1-\alpha)-\phi(z_{\alpha})z_{\alpha}\frac{\sigma_{\dagger,a}-\sigma_{\#,a}}{\sigma_{\#,a}}\right|$
@@
$\quad+{\mathrm{pr}}(|\sqrt{n}\{\tilde{\psi}_{a}-\psi_{*,a}-P_{n}D_{a}(Q_{\#,a},\psi_{*,a})\}|>\tilde{\eta},\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}\leq 2\delta M)$
$\quad+{\mathrm{pr}}\left(|z_{\alpha}(\tilde{\sigma}_{a}-\sigma_{\dagger,a})|>\sum_{j=1}^{8}\tilde{\eta}_{j},\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}\leq 2\delta M\right)+{\mathrm{pr}}(\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}>2\delta M).$
@@
${\mathrm{pr}}(|\sqrt{n}\{\tilde{\psi}_{a}-\psi_{*,a}-P_{n}D_{a}(Q_{\#,a},\psi_{*,a})\}|>\tilde{\eta},\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}\leq 2\delta M)$
$\leq{\mathrm{pr}}(|{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})|>\tilde{\eta},\|H_{a}(\tilde{Q}_{a})\|_{2}\leq\delta F^{H_{a}})\leq\tilde{t}^{-\tilde{q}/2}.$
Next bound ${\mathrm{pr}}\left(|z_{\alpha}(\tilde{\sigma}_{a}-\sigma_{\dagger,a})|>\sum_{j=1}^{8}\tilde{\eta}_{j},\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}\leq 2\delta M\right)$. This term is zero if $z_{\alpha}=0$ and $\sum_{j=1}^{8}\tilde{\eta}_{j}>0$. Otherwise, since $\tilde{\sigma}_{a}\geq 0$ and $\sigma_{\dagger,a}>0$, it holds that
${\mathrm{pr}}\left(|z_{\alpha}(\tilde{\sigma}_{a}-\sigma_{\dagger,a})|>\sum_{j=1}^{8}\tilde{\eta}_{j},\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}\leq 2\delta M\right)$
$\leq{\mathrm{pr}}\left(|\tilde{\sigma}_{a}^{2}-\sigma_{\dagger,a}^{2}|>\sigma_{\dagger,a}\sum_{j=1}^{8}\tilde{\eta}_{j}/|z_{\alpha}|,\|{\mathcal{T}}_{a}(\tilde{Q}_{a})^{2}-{\mathcal{T}}_{a}(Q_{\#,a})^{2}\|_{2}\leq\delta F^{{\mathcal{T}}_{a}^{2}}\right).$
@@
${\mathrm{pr}}(\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}>2\delta M)\leq\frac{{E}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}}{4\delta^{2}M^{2}}.$ (S10)
Additionally under Condition 11,
${\mathrm{pr}}(\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}>2\delta M)\leq\tilde{a}_{1}n^{\tilde{d}_{1}}\exp\left\{-\tilde{c}_{1}\left(\frac{2\delta M}{\tilde{r}_{1}(n)}\right)^{\tilde{q}_{1}}\right\}+\tilde{b}_{1}n^{-1/2}\log n.$ (S11)
@@
$\left|{\mathrm{pr}}(\sqrt{n}(\tilde{\psi}_{a}-\psi_{*,a})\leq z_{\alpha}\tilde{\sigma}_{a})-(1-\alpha)-\phi(z_{\alpha})z_{\alpha}\frac{\sigma_{\dagger,a}-\sigma_{\#,a}}{\sigma_{\#,a}}\right|$
@@
$\quad+{\mathcal{C}}\frac{\varrho_{\#,a}}{\varsigma_{\#,a}^{3}\sqrt{n}}+2\exp\left\{-\frac{\eta_{7}^{2}n}{2\varsigma_{\#,a}^{2}}\right\}+t_{2}^{-q_{2}/2}+{\mathbbm{1}}(\psi_{*,a}\neq 0)t_{4}^{-q_{4}/2}+t_{6}^{-q_{6}/2}+\frac{16M^{2}(1-\tau_{\pi})^{2}{E}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}}{\tau_{\pi}^{2}\eta_{8}^{2}}\Bigg]$
$\quad+\frac{{E}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}}{4\delta^{2}M^{2}},$ (S12)
@@
$\frac{16M^{2}(1-\tau_{\pi})^{2}{E}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}}{\tau_{\pi}^{2}\eta_{8}^{2}}\text{ can be replaced by }\tilde{a}_{2}n^{\tilde{d}_{2}}\exp\left\{-\tilde{c}_{2}\left(\frac{\eta_{8}}{\tilde{r}_{2}(n)}\right)^{\tilde{q}_{2}}\right\}+\tilde{b}_{2}n^{-1/2}\log n$
under Condition 12, and
$\frac{{E}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}}{4\delta^{2}M^{2}}\text{ can be replaced by }\tilde{a}_{1}n^{\tilde{d}_{1}}\exp\left\{-\tilde{c}_{1}\left(\frac{2\delta M}{\tilde{r}_{1}(n)}\right)^{\tilde{q}_{1}}\right\}+\tilde{b}_{1}n^{-1/2}\log n$
under Condition 11.
@@
$J(2\delta,{\mathcal{F}},M)$ $\leq\int_{0}^{2\delta}\sqrt{C_{1}(\nu)+2\nu\log(1/\epsilon)}{\mathrm{d}}\epsilon$